The letters of the alphabet are writtten on slips of paper and placed in a hat. three letters are selected at random.
A. How many different combinations of three letters are possible?
B. How many combinations consist only of the letters A,C,H,I,K or Y?
C. What is the probability that the letters selected consist only of the letters A,C,H,I,K or Y?

Answer by dabanfield(803)

The letters of the alphabet are writtten on slips of paper and placed in a hat. three letters are selected at random.
A. How many different combinations of three letters are possible?
B. How many combinations consist only of the letters A,C,H,I,K or Y?
C. What is the probability that the letters selected consist only of the letters A,C,H,I,K or Y?
A.) The number of ways to select three letters from 26 is
C(26,3) = 26!/((26-3)!*3!) = 26!/(23!*3!) = (26*25*24)/(3*2*1) = ?
B.) The number of ways to select three letters from the group of 6 letters is
C(6,3) = 6!/((6-3)!*3!) = 6!/(3!*3!) = (6*5*4)/(3*2*1) = ?
C.) (answer from A)/(answer from B)